OpenStudy (anonymous):

Given det(A) = -7 and det(B) = 8, what is the determinant of (5A^-1)*(B^T)

OpenStudy (anonymous):

I know that the det(B^T) is the same and the det(A^-1) = 1/det(A) so I did 5*(1/-7)*8 = 40/-7 which was not correct :c

OpenStudy (amistre64):

B^T is fine 2 3 4 5 det = 10 - 12 = -2 inverse 5/-2 -3/-2 -4/-2 2/-2 10/4 - 12/4 = -2/4 = -1/2 the rules seem to hold

OpenStudy (amistre64):

5k/-2 -3k/-2 -4k/-2 2k/-2 k*k = k^2 to

OpenStudy (amistre64):

det(kA) = k^2 det(A) is what i see it at

OpenStudy (amistre64):

but im not up to date on the rules

OpenStudy (anonymous):

Assuming \(A\) and \(B\) are square with the same dimensions, you have \[\det (AB)=\det A\times\det B\] which must be the case, otherwise \(A\) wouldn't have an inverse and \(A^{-1}B^T\) wouldn't exist.